OpenStudy (anonymous):

3. What is a polynomial function in standard form with zeroes 1, 2, –3, and –3? (1 point) A. g(x) = x4 + 3x3 –7x2 – 15x + 18 B. g(x) = x4 + 3x3 –7x2 + 2x + 18 C. g(x) = x4 – 3x3 + 7x2 + 15x + 18 D. g(x) = x4 – 3x3 –7x2 + 15x + 18

OpenStudy (hba):

Jus plug them in @kimblack

OpenStudy (anonymous):

just plug them in for x ?

OpenStudy (hba):

Yo black.

OpenStudy (anonymous):

can you do one for example ? ... letter a

OpenStudy (hba):

Nah,I don't have time for that.

OpenStudy (anonymous):

did you get this answer @kimblack?

OpenStudy (anonymous):

it is A (:

OpenStudy (anonymous):

oh my gosh, you are the best. thank you!

OpenStudy (anonymous):

yupp.. do you have anymore questions on that test ???

OpenStudy (anonymous):

haha umm, I have no idea what I'm doing on any part of this test so any help would be suuuper.

OpenStudy (anonymous):

well... tell me what answers you have.. and i'll give you the rest.

OpenStudy (anonymous):

1,2,3,6,8,10,11 aaand that's it.

OpenStudy (anonymous):

4.d 5.b 7.a 9.b 12.d 13.d 14. This is your equation.. simplify 1.5x+x=6 16.151 18.1d^6-30d^5+375d^4-2500d^3+9375d^2-18750d+15625 19.y=-1/7x^3(x+8) .. you will only get partial credit for this one the other one's you got look up.. or ask someone.. soo (: there you go

OpenStudy (anonymous):

Thanks so much! :) Honestly you have no idea, how much this helps.

OpenStudy (anonymous):

1.b 2.b 3.a 6.c 8.b 10.a 11.b here are those just incase you got those wrong..

OpenStudy (anonymous):

Finally just finished doing them all. You're a savior, thank you again.

OpenStudy (anonymous):

you're welcome.. what did you get on it ?